# A Unified Multi-Wavelength Data Analysis Workflow with gammapy

### v1

@@ 3.2 NuSTAR data reduction
The Nuclear Spectroscopic Telescope Array (NuSTAR, Harrison et al. 2013) is a focusing high-energy X-ray telescope operating in the 3 to 79 keV energy range. NuSTAR is a telescope array consisting on two twin modules (NuSTAR A and NuSTAR B), each made of four Cadmium-Zinc-Tellurium detectors.
The standard data products provided to the observer are similar to other X-ray space telescopes and consist of an event list where the most important columns are the reconstructed position, energy and time of arrival of each detected X-ray photon.
@@
In a second step, we ran nuproducts (Stage 3) over the Level 2 event files to extract the integrated bandpass images ($3-79\,$keV), the source spectrum and light curves. Each NuSTAR module and sector has slightly different properties, therefore the background extraction regions were optimized for each module in shape and location to maximize background statistics, but keeping it within the limits of the same sector in which the source is located (see Fig. 3). From this point, we focus only on the spectral analysis of the Level 3 data produced by nuproducts, consisting of: (i) PHA files (energy spectra) for the source and background; (ii) ARF file (Ancilliary Response File), which contains the effective area; and (iii) RMF file (Response or Redistribution Matrix File), containing the energy migration matrix, a 2D-array that gives the probabilities of assigning the photon to a given PHA channel for each input photon energy.
@@ 3.2.1 Background model
Two analyses of the spectral data were considered, each with different background assumptions. The first analysis uses the ON-OFF method with W-statistics (W-statistics, wstat, or Poisson data with a Poisson background). In this approach, the background is estimated measuring the number of events in the OFF region (see Fig.3), assuming that it is the same as in the source region once correcting for the region size differences. This method is straightforward but can not account for any spatial variation of the background (e.g. instrumental lines). In addition, it is statistically limited, as the W-statistic estimator can produce inaccurate results in low-count regimes, particularly if the background reaches zero counts in a given energy bin. To mitigate these issues, we re-binned the data into coarser energy bins to increase the statistics per bin.
The second approach involves using a background model for NuSTAR data. We employed the Cosmic X-ray Background (CXB) simulation tool nuskybgd (Wik et al. 2014) to estimate the background (using C-statistics, cstat, or Poisson data with a background model). Compared to a directly measured background via the OFF background method, the predicted background from nuskybgd effectively addresses issues related to very low statistics at high energies. At the cost of the complexity added by nuskybgd, this approach retains better spectral resolution while fully utilizing the energy range of NuSTAR, and accounts for the spatial variation of instrumental background. A detailed comparison of the impact of the choice of the background method is presented in Appendix B.
@@ 3.2.2 Dataset and IRFs
Figure 4: gammapy dataset representation of NuSTAR A for night 1 is shown here. NuSTAR B for this observation and NuSTAR A/B for night 2 are similar. From top left to bottom right: (a) Estimated distribution of background events from the OFF rectangular region of Figure 3 and ON event distribution from the signal extraction region. (b) Estimated distribution of background events from nuskybgd. ON event distribution identical to the previous case. (c) NuSTAR A exposure (effective area $\times$ livetime) as a function of true energy. (d) NuSTAR A energy dispersion matrix.
@@ 3.3 Swift-XRT data reduction
The X-Ray Telescope (XRT, Burrows et al. 2004) onboard the Swift satellite is an X-ray CCD imaging spectrometer designed to measure the position, spectrum, and brightness of astronomical objects in the range of $0.2$ to $\sim 10\,$ keV, with a dynamic range spanning more than seven orders of magnitude in flux. XRT has an effective area of approximately $135\,\mathrm{cm^{2}}$ at 1.5 keV and a detection sensitivity of $2\times 10^{-14}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ in 101010\,10 ks of observing time.
The FoV is $23.6\times 23.6$ arcmin, with an angular resolution of $18$ arcsec (Half-Energy Width, HEW). XRT operates in two readout modes: Windowed Timing (WT) and Photon Counting (PC). The WT mode bins 10 rows into a single row, focusing on the central 200 columns to improve time resolution to 1.71.71.7\,1.7 ms (ideal for bright sources), while the PC mode retains full imaging and spectroscopic capabilities but with a slower time resolution of $2.5$ s and is limited by pileup effects at high count rates ($\gtrsim 0.6\,$ c/s, Vaughan et al. 2006). OP 313 is a relatively weak X-ray source with a modest count rate of $\lesssim 0.2\,$ c/s, well within the capabilities of XRT in PC mode; therefore, this work focuses exclusively on this mode.
Swift-XRT data for OP 313 were downloaded from the UK Swift Science Data Centre [^4] https://www.swift.ac.uk/swift_portal/. Of all available data from the telescope, this work focuses on data collected during the nights of March 4th, 2024 and March 15th, 2024, with Swift observation IDs 00036384074 (2.962.962.96\,2.96ks) and 00089816002 (3.163.163.16\,3.16ks), respectively. We used Heasoft 6.32.1 and xrtdas v3.7.0 for data reduction. Spectra were extracted and saved using ximage for fixed source and background regions (see Fig. 5) in two separate PHA files. The corresponding Response Matrix File (RMF) file was downloaded from HEASARC’s calibration database (CALDB), and the Ancilliary Response File (ARF) was produced using xrtmkarf.
@@ 3.3.1 Dataset and IRFs
The source and background PHA files, together with the RMF and ARF, were used to generate gammapy’s SpectrumDatasetOnOff objects, represented in Fig. 5. Following the discussion for NuSTAR on background statistics, we re-binned the spectral data into coarser energy bins, aiming at avoiding zero counts in the background bins and therefore preserving the accuracy of the W-statistics assumption.
Figure 5: Swift-XRT dataset representation after energy bin grouping to avoid a background spectra dominated by zeros. (a) Swift-XRT image of OP 313. Blue circle and orange annulus show the signal and background extraction regions, respectively. (b) Distribution of events for both regions as a function of energy, corrected for the different region sizes. (c) Swift-XRT exposure (effective area $\times$ livetime) as a function of true energy. (d) Swift-XRT energy dispersion matrix (also known as redistribution matrix or migration matrix).
@@ 3.4 Swift-UVOT data reduction
The Ultraviolet/Optical Telescope (UVOT, Roming et al. 2005) onboard the Swift satellite is a diffraction-limited 303030\,30cm modified Ritchey-Chrétien telescope with a limiting magnitude of $22.3$ in 111\,1 ks exposures. The CCD detector, based on XMM-Newton’s CCD (MIC) detectors, operates in photon counting mode over the range of $170$ to 650650650\,650 nm with 7 filters (V, B, U in the visible band; UVW1, UVM2, UVW2 in the UV band; and a white or full-passband filter).
Data reduction for Swift-UVOT is performed using heasoft v6.32.1 and the Python package photutils (Bradley et al. 2022). After downloading the target images from HEASARC’s repository, we stack images from different filters, aligning them using their world coordinate system (WCS) solutions to maximize the signal-to-noise ratio. This approach helps define the source and background regions accurately, even for dim targets, and allows to better mask potential sources in the field that might bias the background extraction. The source extraction region is defined as a circular aperture with a radius of 555\,5 arcsec, initially centered on the nominal position of OP 313 and adjusted based on centroid calculations on the stacked image. A background extraction region is then defined as an annulus with an inner radius of 12.512.512.5\,12.5 arcsec and an outer radius of 404040\,40 arcsec, centered on the source region. Bright point-like sources in the FoV, identified by photutils, are accounted for by adding small ’masking’ circles to the background region file. An illustration of this method, including a cutout of the Swift-UVOT images for different filters, is shown in Figure 6. OP 313 is a distant, strongly Doppler-boosted source that appears point-like in most energy bands, including optical and UV. Consequently, UVOT data are analyzed using the same strategy used for XRT and NuSTAR. The result is the production of PHA source and background files (containing the integrated number of counts, single channel) for each filter. UVOT Response files (RSP) are downloaded directly from CALDB.
@@ 3.4.1 Dataset and IRFs
In imaging mode (i.e., without using the grisms) and for point-like sources, we treat UVOT as a single-channel detector with a response determined by the selected photometric filter. The RSP files provide effective area information for each filter (see Figure 6, second row). A dummy redistribution matrix is generated on the fly to match gammapy’s expected format definition. These redistribution matrices consist of a row of constant values, indicating that all photon energies contribute to the single channel of the detector.
Figure 6: First row: Swift-UVOT images in two of the photometric filters available for the first selected night, with signal and background extraction regions indicated by blue and orange dashed circles and annuli, respectively. Second row: Effective collection area for each of the available Swift-UVOT filters.
@@ 3.5 Liverpool-IO:O data reduction
We developed a custom data analysis pipeline using the photutils Python package to process LT IO:O calibrated images through differential aperture photometry. This pipeline reduces each image into a SpectrumDatasetOnOff compatible format in gammapy, calculating signal and background counts in ON and OFF regions, respectively. The exposure or instrument response is estimated with the same images, using a differential photometry technique consisting on measuring the flux of several calibration stars in the FoV and comparing it to their known flux densities in reference catalogs.
@@ 3.5.2 Exposure estimation
For optical ground-based instruments, characterizing $\mathcal{E}(E_{\rm true})$ is complex due to the need to account for both optical elements and atmospheric effects, the latter particularly hard to assess and only accounted for as an energy-averaged scaling factor of the exposure which is derived directly from star photometry. We approximate the energy-dependence of the exposure by the detector throughput curves provided in Barnsley et al. (2016), consisting of the product of filter transmission times cryostat transmission times CCD quantum efficiency (see Figure 7). This approximation assumes that the photometric band is narrow enough for the source spectrum to be approximated by the photon density $F_{\nu}/\nu_{c}$, where $\nu_{c}$ is the effective frequency of the photometric band and $F_{\nu}$ is the flux density. Rewriting Eq. 3 in terms of wavelength, we get:
@@
¯$\langle\bar{\mathcal{E}}\rangle=\frac{\sum_{k}\frac{\bar{\mathcal{E}}_{k}}{\sigma_{\bar{\mathcal{E}}_{k}}^{2}}}{\sum_{k}\frac{1}{\sigma_{\bar{\mathcal{E}}_{k}}^{2}}}\pm\sqrt{\frac{1}{\sum_{k}\frac{1}{\sigma_{\bar{\mathcal{E}}_{k}}^{2}}}+\sum_{k}(\bar{\mathcal{E}}_{k}-\langle\bar{\mathcal{E}}\rangle)^{2}}.$ (5)
@@ 3.5.3 Dataset and IRFs
Figure 7: Top row: LT IO:O images in two of the photometric filters available for the first selected night, with the signal and background extraction regions represented by blue and orange dashed circles and annuli, respectively. Bottom row: Throughput for each of the available LT IO:O bands (transmission of the filter multiplied by the cryostat window transmission, CW, and the quantum efficiency of the CCD, QE), used to shape the energy dependence of the exposure in LT. The curves for the quantum efficiency multiplied by the cryostat, and for the cryostat alone are shown for completeness.
@@
In the forward-folding approach described here, an alternative method is utilized to include normalization models for each dataset or each filter (e.g., using a PiecewiseNormSpectralModel with normalization factors defined at filter-center energies). Gaussian priors are applied to account for the measured systematic uncertainties, with the widths of the Gaussian priors set to the standard deviation of the relative errors in the exposure with respect to the calculated mean. These relative errors range from less than 0.10.10.1\,0.1mag for the $z$- and $i$-bands to more than 0.30.30.3\,0.3mag for the $u$-band, where fewer reference stars are detectable. A more detailed discussion on the practical differences between the two methods and the effect of adding the systematic in quadrature to the statistical uncertainty is provided in Appendix D.
@@ 4.1 Summary of observational data and analysis
We present the results from the broadband emission analysis of OP 313, spanning from optical to high-energy $\gamma$ rays. The instruments used, listed in decreasing order of energy coverage, are: Fermi-LAT, NuSTAR, Swift-XRT, Swift-UVOT, and the LT with the IO:O camera. Our analysis aims to interpret the data in a multi-wavelength context and evaluate the capabilities of our new multi-wavelength data management workflow. Before testing broadband modeling of the joint dataset, we carried out a technical validation of the analysis of data for each instrument separately, comparing in each case the spectral reconstruction of gammapy against that of the analysis tools native to each instrument. The details of this validation can be found in Appendix C. This section is divided into three parts:
@@
Figure 8: Temporal coverage of the first observation night with the different participating instruments.
@@ 4.2.1 Emission components
To model the broadband emission from OP 313, we employed a phenomenological approach using two power-laws with exponential cut-offs with index $\alpha=1$ fixed:

### v2

@@ 3.2 NuSTAR data reduction
The Nuclear Spectroscopic Telescope Array (NuSTAR, Harrison et al. 2013) is a focusing high-energy X-ray telescope operating in the 3 to 79 keV energy range. NuSTAR is a telescope array consisting of two twin modules (NuSTAR A and NuSTAR B), each made of four Cadmium-Zinc-Tellurium detectors.
The standard data products provided to the observer are similar to other X-ray and gamma-ray space telescopes and consist of an event list with the reconstructed position, energy and time of arrival of arrival of each detected photon.
@@
In a second step, we ran nuproducts (Stage 3) over the Level 2 event files to extract the integrated bandpass images ($3-79\,$keV), the source spectrum and light curves. Each NuSTAR module and sector has slightly different properties, therefore the background extraction regions were optimized for each module in shape and location to maximize background statistics, but keeping it within the limits of the same sector in which the source is located (see Fig. 3). From this point, we focus only on the spectral analysis of the Level 3 data produced by nuproducts, consisting of: (i) PHA files (energy spectra) for the source and background; (ii) ARF file (Ancillary Response File), which contains the effective area; and (iii) RMF file (Response or Redistribution Matrix File), containing the energy migration matrix, a 2D-array that gives the probabilities of assigning the photon to a given PHA channel for each input photon energy.
@@ 3.2.1 Background model
Two analyses of the spectral data were considered, each with different background assumptions. The first analysis uses the ON-OFF method with W-statistics (W-statistics, wstat [^5] Note that in xspec, no name distinctions are made between cstat and wstat, see https://heasarc.gsfc.nasa.gov/docs/xanadu/xspec/manual/XSappendixStatistics.html, or Poisson data with a Poisson background). In this approach, the background is estimated measuring the number of events in the OFF region (see Fig.3), assuming that it is the same as in the source region once correcting for the region size differences. This method is straightforward but cannot account for any spatial variation of the background (e.g. instrumental lines). In addition, it is statistically limited, as the W-statistic estimator can produce inaccurate results in low-count regimes, particularly if the background reaches zero counts in a given energy bin. To mitigate these issues, we re-binned the data into coarser energy bins (by a factor of 32) to increase the statistics per bin.
The second approach involves using a background model for NuSTAR data. We employed the Cosmic X-ray Background (CXB) simulation tool nuskybgd (Wik et al. 2014) to estimate the background (using C-statistics, cstat, or Poisson data with a background model). Compared to a directly measured background via the OFF background method, the predicted background from nuskybgd effectively addresses issues related to very low statistics at high energies. At the cost of the complexity added by nuskybgd, this approach retains better spectral resolution while fully utilizing the energy range of NuSTAR, and accounts for the spatial variation of instrumental background. A detailed comparison of the impact of the choice of the background method is presented in Appendix D.
@@ 3.2.2 Dataset and IRFs
Figure 4: gammapy dataset representation of NuSTAR A for night 1 is shown here. NuSTAR B for this observation and NuSTAR A/B for night 2 are similar. From top left to bottom right: (a) Estimated distribution of background events from the OFF rectangular region of Figure 3 and the ON event distribution from the signal extraction region. (b) Estimated distribution of background events from nuskybgd. ON event distribution identical to the previous case. (c) NuSTAR A exposure (effective area $\times$ livetime) as a function of true energy. (d) NuSTAR A energy dispersion matrix.
@@ 3.3 Swift-XRT data reduction
The X-Ray Telescope (XRT, Burrows et al. 2004) onboard the Swift satellite is an X-ray CCD imaging spectrometer designed to measure the position, spectrum, and brightness of astronomical objects in the range of $0.2,$ to $10,$ keV, with a dynamic range spanning more than seven orders of magnitude in flux. XRT has an effective area of approximately $135\,\mathrm{cm^{2}}$ at 1.5 keV and a detection sensitivity of $2\times 10^{-14}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$ in 101010\,10 ks of observing time.
The FoV is $23.6\times 23.6$ arcmin, with an angular resolution of $18$ arcsec (Half-Energy Width, HEW). XRT operates in two readout modes: Windowed Timing (WT) and Photon Counting (PC). The WT mode bins 10 rows into a single row, focusing on the central 200 columns to improve time resolution to 1.71.71.7\,1.7 ms (ideal for bright sources). The PC mode retains full imaging and spectroscopic capabilities but with a slower time resolution of $2.5$ s, and is limited by pileup effects at high count rates ($\gtrsim 0.6\,$ c/s, Vaughan et al. 2006). OP 313 is a relatively weak X-ray source with a modest count rate of $\lesssim 0.2\,$ c/s, well within the capabilities of XRT in PC mode; therefore, this work focuses exclusively on this mode.
3.3.1 Data reduction
Swift-XRT data for OP 313 were downloaded from the UK Swift Science Data Centre [^6] https://www.swift.ac.uk/swift_portal/. Of all available data from the telescope, this work focuses on data collected during the nights of March 4th, 2024 and March 15th, 2024, with Swift observation IDs 00036384074 (2.962.962.96\,2.96ks) and 00089816002 (3.163.163.16\,3.16ks), respectively. We used Heasoft 6.32.1 and xrtdas v3.7.0 for data reduction. Spectra were extracted using xrtproducts for a fixed source extraction region of $40"$ radius, and a surrounding annular region of $80"$ and $300"$ radii for the background (see Fig. 5). This process results in two separate PHA files, along with the corresponding ARF. The RMF file was downloaded from HEASARC’s calibration database (CALDB).
@@ 3.3.2 Dataset and IRFs
The source and background PHA files, together with the RMF and ARF, were used to generate gammapy’s SpectrumDatasetOnOff objects, represented in Fig. 5. Following the discussion for NuSTAR on background statistics, we re-binned the spectral data into coarser energy bins with 30 channels each, to avoid zero counts in the background bins and therefore preserving the accuracy of the W-statistics assumption.
Figure 5: Swift-XRT dataset representation after energy bin grouping to avoid background spectra dominated by zeros. (a) Swift-XRT image of OP 313. Blue circle and orange annulus show the signal and background extraction regions, respectively. (b) Distribution of events for both regions as a function of energy, corrected for the different region sizes. (c) Swift-XRT exposure (effective area $\times$ livetime) as a function of true energy. (d) Swift-XRT energy dispersion matrix (also known as redistribution matrix or migration matrix).
@@ 3.4 Swift-UVOT data reduction
The Ultraviolet/Optical Telescope (UVOT, Roming et al. 2005) onboard the Swift satellite is a diffraction-limited 303030\,30cm modified Ritchey-Chrétien telescope with a limiting magnitude of $22.3$ in 111\,1 ks exposures. The UVOT detector, identical to XMM-Newton’s microchannel-plate intensified CCD, operates in photon counting mode over the range of $170$ to 650650650\,650 nm with 7 filters (V, B, U in the visible band; UVW1, UVM2, UVW2 in the UV band; and a white or full-passband filter).
3.4.1 Data reduction
Data reduction for Swift-UVOT is performed using HEASoft v6.32.1. After downloading the target images from HEASARC’s repository, source extraction regions were defined as circular apertures. We used an exceptionally large radius of $25"$ as one Swift gyroscope gradually degraded since July 2023 until failing in March 2024, increasing the likelihood of star tracker “loss of lock” after long slews (Cenko 2024a, b). The effect was particularly strong in the second night. The source regions are initially centered on the nominal position of OP 313 and fine-tuned based on centroid calculations. A background extraction region is then defined as an annulus with an inner radius of $40"$ and an outer radius of $100"$, centered on the source region. An example of the source and background regions, overlaid on top of the cutout of the Swift-UVOT images for different filters, is shown in Figure 6. OP 313 is a distant, strongly Doppler-boosted source that appears point-like in most energy bands, including optical and UV. Consequently, UVOT data are analyzed using the same strategy used for XRT and NuSTAR. The result is the production of PHA source and background files (containing the integrated number of counts) for each filter. UVOT Response files (RSP) are downloaded directly from CALDB.
@@ 3.4.2 Dataset and IRFs
In imaging mode (i.e., without using the grisms) and for point-like sources, we treat UVOT as a single-channel detector with a response determined by the selected photometric filter. The RSP files provide effective area information for each filter (see Figure 6, second row). A dummy redistribution matrix is generated on the fly to match gammapy’s expected format definition. These redistribution matrices consist of a row of constant values, indicating that all photon energies contribute to the single-channel detector.
Figure 6: First row: Swift-UVOT images in two of the photometric filters available for the first selected night, indicated by dashed blue circles (signal) and orange annuli (background), respectively. Second row: Effective collection area for each of the available Swift-UVOT filters.
@@ 3.5 Liverpool-IO:O data reduction
We developed a custom data analysis pipeline using the photutils(Bradley et al. 2022) Python package to process LT IO:O calibrated images through differential aperture photometry. This pipeline reduces each image into a SpectrumDatasetOnOff compatible format in gammapy, calculating signal and background counts in ON and OFF regions, respectively. The exposure or instrument response is estimated with the same images, using a differential photometry technique consisting on measuring the flux of several calibration stars in the FoV and comparing it to their known flux densities in reference catalogs.
@@ 3.5.2 Exposure estimation
For optical ground-based instruments, characterizing $\mathcal{E}(E_{\rm true})$ is complex due to the need to account for both optical elements and atmospheric effects, the latter particularly hard to assess and only accounted for as an energy-averaged scaling factor of the exposure which is derived directly from star photometry. We approximate the energy-dependence of the exposure (effective area multiplied by livetime, as per gammapy definition) by the detector throughput curves provided in Barnsley et al. (2016), consisting of the product of filter transmission times cryostat transmission times CCD quantum efficiency (see Figure 7). This approximation assumes that the photometric band is narrow enough for the source spectrum to be approximated by the photon density $F_{\nu}/\nu_{c}$, where $\nu_{c}$ is the effective frequency of the photometric band and $F_{\nu}$ is the flux density. Rewriting Eq. 3 in terms of wavelength, we get:
@@
¯$\langle\bar{\mathcal{E}}\rangle=\frac{\sum_{k}\frac{\bar{\mathcal{E}}_{k}}{\sigma_{\bar{\mathcal{E}}_{k}}^{2}}}{\sum_{k}\frac{1}{\sigma_{\bar{\mathcal{E}}_{k}}^{2}}},\\ \Delta\langle\bar{\mathcal{E}}\rangle=\sqrt{\frac{1}{\sum_{k}\frac{1}{\sigma_{\bar{\mathcal{E}}_{k}}^{2}}}+\sum_{k}(\bar{\mathcal{E}}_{k}-\langle\bar{\mathcal{E}}\rangle)^{2}}.$ (5)
@@ 3.5.3 Dataset and IRFs
Figure 7: Top row: LT IO:O images in two of the photometric filters available for the first selected night, with the signal and background extraction regions represented by blue and orange dashed circles and annuli, respectively. Bottom row: Throughput for each of the available LT IO:O bands (transmission of the filter multiplied by the cryostat window transmission, CW, and the quantum efficiency of the CCD, QE) as dashed curves, and for CW and CW×$\times$×QE alone. For completeness, the reconstructed effective area for each calibration star and each filter are shown as solid colored curves.
@@
In the forward-folding approach described here, an alternative method is utilized to include normalization models for each dataset or each filter (e.g., using a PiecewiseNormSpectralModel with normalization factors defined at filter-center energies). Gaussian priors are applied to account for the measured systematic uncertainties, with the widths of the Gaussian priors set to the standard deviation of the relative errors in the exposure with respect to the calculated mean. These relative errors range from less than 0.10.10.1\,0.1mag for the $z$- and $i$-bands to more than 0.30.30.3\,0.3mag for the $u$-band, where fewer reference stars are detectable. A more detailed discussion on the practical differences between the two methods and the effect of adding the systematic in quadrature to the statistical uncertainty is provided in Appendix E.
@@ 4.1 Summary of observational data and analysis
We present the results from the broadband emission analysis of OP 313, spanning from optical to high-energy $\gamma$ rays. The instruments used, listed in decreasing order of energy coverage, are: Fermi-LAT, NuSTAR, Swift-XRT, Swift-UVOT, and the LT with the IO:O camera. Our analysis aims to interpret the data in a multi-wavelength context and evaluate the capabilities of our new multi-instrument data management workflow. Before testing broadband modeling of the joint dataset, we carried out a technical validation of the analysis of data for each instrument separately, comparing in each case the spectral reconstruction of gammapy against that of the analysis tools native to each instrument. The details of this validation can be found in Appendix C. This section is divided into three parts:
@@
Figure 8: Temporal coverage of the first observation night.
@@ 4.2.1 Emission components
To model the broadband emission from OP 313, we employed a phenomenological approach using two power-laws with exponential cut-offs with a fixed index of $\alpha=1$: